OpenStudy (anonymous):

I need help on solving exponential equations, because I wasn't in class yesterday while we learned it. Please help me!? 25^2x+3=25^5x-9

jimthompson5910 (jim_thompson5910):

The equation is this right? \[\Large 25^{2x+3} = 25^{5x-9}\]

jimthompson5910 (jim_thompson5910):

If so, then we can equate the exponents. This is because the bases are equal, so in order for both sides to be equal, the exponents must be equal that means \[\Large 25^{2x+3} = 25^{5x-9}\] leads to \[\Large 2x+3 = 5x-9\] solve for x to get your answer

jimthompson5910 (jim_thompson5910):

this is assuming my initial assumption about the problem is correct

OpenStudy (anonymous):

Yes that is the right equation I typed it on my phone so it was weird, and thank you so much! That helped a lot!

jimthompson5910 (jim_thompson5910):

I'm glad it makes more sense now

OpenStudy (anonymous):

Okay so now can you please help me on how to get the bases equal? please? \[9^{8x-4}=81^{3x+6}\]

jimthompson5910 (jim_thompson5910):

the bases aren't equal, so we can't equate the exponents like we did in the prev problem but there's a simple fix here notice how 81 is equal to 9*9 or 9^2 so we can do this bit of algebraic manipulation \[\Large 9^{8x-4} = 81^{3x+6}\] \[\Large 9^{8x-4} = \color{red}{81}^{3x+6}\] \[\Large 9^{8x-4} = (\color{red}{81})^{3x+6}\] \[\Large 9^{8x-4} = \left(\color{red}{9^2}\right)^{3x+6}\] \[\Large 9^{8x-4} = \left(9^2\right)^{3x+6}\] \[\Large 9^{8x-4} = 9^{2(3x+6)}\]

jimthompson5910 (jim_thompson5910):

On that last step, I used the rule \[\Large \left(a^b\right)^c = a^{b*c}\]

jimthompson5910 (jim_thompson5910):

now that the bases are both equal to 9, you can now equate the exponents and solve for x

OpenStudy (anonymous):

Thank you so much. Definitely understand it better now.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Can you help me work through solving a logarithm equation? Please. \[1/2\log6(25)+\log6{x} =\log6{20} \]

jimthompson5910 (jim_thompson5910):

the '6' is the base of each log?